Evaluate the integral "\oint" ezt/z2+1 dz where c is the circle |z|=3.



1
Expert's answer
2021-09-10T15:17:16-0400

Inside the contour "C=\{z\in\mathbb{C}:|z|=3\}", the function "\frac{e^{zt}}{z^2+1}" has 2 simple poles at "z=i" and "z=-i". Hence, according to the residue theorem,

"\oint\limits_C \frac{e^{zt}}{z^2+1}dz=2\pi i({\rm res}_{z=i}\frac{e^{zt}}{z^2+1}+{\rm res}_{z=-i}\frac{e^{zt}}{z^2+1})"

"\operatorname{res}_{z=i}\frac{e^{zt}}{z^2+1}=\lim\limits_{z\to i}\frac{e^{zt}}{z^2+1}(z-i)=\lim\limits_{z\to i}\frac{e^{zt}}{z
+i}=\frac{e^{it}}{2i}"

"{\rm res}_{z=-i}\frac{e^{zt}}{z^2+1}=\lim\limits_{z\to i}\frac{e^{zt}}{z^2+1}(z+i)=\lim\limits_{z\to i}\frac{e^{zt}}{z
-i}=-\frac{e^{-it}}{2i}"

Therefore,

"\oint\limits_C \frac{e^{zt}}{z^2+1}dz=2\pi i(\frac{e^{it}}{2i}-\frac{e^{-it}}{2i})=2\pi i\sin t"


Answer. "\oint\limits_C \frac{e^{zt}}{z^2+1}dz=2\pi i\sin t"